# Srinivasan Kidambi, and Pravin Nair

### v1

@@ Abstract
Index Terms:
Attention, Image and video restoration, Training-free, Plug-and-Play, Nyström approximation.
I Introduction
Transformers have recently demonstrated strong performance across a range of low-level vision tasks, including image denoising, deblurring [1], super-resolution [2], and video restoration [3, 4]. These models rely on a self-attention mechanism that captures long-range dependencies by computing pairwise similarity between features corresponding to image or video patches. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{i}}$ be the input feature map corresponding to $N$ tokens (e.g., pixels or patches), where each token is represented by a $d_{i}$-dimensional feature vector. To compute self-attention [5], the input $\mathbf{X}$ is first projected into query, key, and value representations:
@@ I Introduction
However, the attention operation in (1) incurs a quadratic complexity in the number of input tokens $N$, which corresponds to the number of pixels or patches within a subregion of a frame [5]. In particular, the time and storage complexity for computing (1) is $\mathcal{O}(N^{2}d)$ and $\mathcal{O}(N^{2}+Nd)$, respectively, dominated by storing and computing the large attention matrices $\mathbf{A}$ and $\mathbf{S}$. Currently, most vision models employ Multi-Head Self-Attention (MHSA) [5], an extension of self-attention. In MHSA, the embedding dimension $d$ is divided into $h$ attention heads, each of dimension $d_{h}$, such that $d=hd_{h}$. For each head, self-attention outputs are independently computed, concatenated, and linearly projected to form MHSA output. The time and storage complexity of MHSA is quadratic in $N$, similar to self-attention. This computational bottleneck makes transformers less suitable for real-time or resource-constrained settings, where convolutional networks are still often preferred due to their computational efficiency [6, 7].
Several approaches approximate self-attention via low-rank [8, 9, 10] or sparse [11] representations to reduce quadratic complexity, but these methods require retraining. Linformer [8] projects keys/values into lower-dimensional subspaces via learned linear maps, achieving linear complexity in $N$. Performer [9] replaces the softmax kernel with random feature approximation, resulting in linear complexity and unbiased attention estimates. Nyströmformer [10] heuristically reconstructs attention using landmark-based sampling. Axial Attention [12] factorizes 2D attention into 1D along spatial axes, reducing cost from $H^{2}W^{2}$ to $HW(H+W)$, where the number of tokens is $N=HW$. FlashAttention [13] and Triton [14] preserve exact attention while optimizing execution through custom fused GPU kernels but are hardware-specific and difficult to integrate into pre-trained models.
@@
1) Linear attention framework: Unlike MHSA, which incurs quadratic time and memory complexity in the number of tokens $N$, PnP-Nystra achieves linear scaling in both time and memory.
2) Significant acceleration: PnP-Nystra can replace standard attention layers in existing window-attention-based transformer models for image and video restoration, yielding substantial speed-ups. For example, in a video super-resolution setting with an upscaling factor of $2$, we observe up to 4×4$\times$4 × speed-up on GPU compared to the original pre-trained model.
3) Robust performance without retraining: Despite significant acceleration, the reduction in restoration quality is minimal. For instance, in image denoising, PnP-Nystra achieves up to 3.5×3.5$\times$3.5 × speed-up on CPU with $<1.3$ dB PSNR reduction.
To the best of our knowledge, this is the first approach to enable inference-time replacement of self-attention in pre-trained transformers without any fine-tuning. This is especially valuable for real-time settings where retraining is impractical.
II Proposed Method
@@
Note that, while the Nyströmformer [10] also employs the Nyström method, it applies the Nyström method directly to the softmax-normalized attention matrix $\mathbf{S}$. Specifically, following the structure in Eq. 3, the authors define:
$\widetilde{\mathbf{S}}=\operatorname{softmax}\left(\frac{\widetilde{\mathbf{Q}}\widetilde{\mathbf{K}}^{\top}}{\sqrt{d}}\right).$
Unlike the kernel matrix $\widetilde{\mathbf{G}}$ in our formulation, $\widetilde{\mathbf{S}}$ cannot be block-partitioned due to row-wise normalization, which induces global coupling among elements. Consequently, a direct Nyström factorization is not applicable. Instead,  [10] proposes a heuristic approximation:
$\mathbf{S}\!\approx\!\operatorname{softmax}\!\left(\frac{\mathbf{Q}\bar{\mathbf{K}}^{\top}}{\sqrt{d}}\right)\!\!\left[\operatorname{softmax}\!\left(\frac{\bar{\mathbf{Q}}\bar{\mathbf{K}}^{\top}}{\sqrt{d}}\right)\right]^{\!\dagger}\!\!\operatorname{softmax}\!\left(\frac{\bar{\mathbf{Q}}\mathbf{K}^{\top}}{\sqrt{d}}\right)$
While computationally efficient, this construction lacks theoretical approximation guarantees such as those in Lemma II.1, in contrast to our PnP-Nystra, which is grounded in provable low-rank kernel approximations.
II-A Computational Details
Standard self-attention incurs a time complexity of $\mathcal{O}(N^{2}d)$ and memory usage of $\mathcal{O}(N^{2}+Nd)$. In contrast, our proposed method reduces both significantly by avoiding full $N\times N$ attention matrix computation. The per-step complexity is summarized in Table I. The overall time complexity is reduced to $\mathcal{O}(Nm(d+d_{v}))$ and memory usage is reduced to $\mathcal{O}(m^{2}+N(m+d_{v}))$ which is significantly lower than standard attention when $m\ll N$.
TABLE I: Per-Step Time and Space Complexity of PnP-Nystra
Complexity Step 2 Step 3 Step 4 Step 5 Step 6 Step 7
@@ II-A Computational Details
In Step $1$ of Algorithm 1, we construct landmark queries and keys by averaging feature tokens over non-overlapping spatial windows [23]. This approach is computationally efficient and well-aligned with the local image statistics. In Step $4$, we observed that the built-in PyTorch pseudoinverse function produces numerically unstable results in practice. To address this, we adopt the robust pseudoinverse computation technique from [24], an iterative algorithm. Additionally, to avoid overflow during kernel computation in Steps 2 and 3, we replace large exponentials beyond a threshold with linear approximations to ensure numerical stability.
II-B Approximation error
Lemma II.1 (Spectral norm error bound)
Assume $\mathbf{G}_{A}$ is nonsingular. Then,
where $\sigma_{k}(\cdot)$ denotes the $k$-th singular value, and the constant $C$ depends quadratically on $\sigma_{1}(\widetilde{\mathbf{G}})$.
The proof follows from Lemma 5.10 in [15], along with the fact that the spectral norm error between $\mathbf{G}$ and $\widehat{\mathbf{G}}$ is upper bounded by that between $\widetilde{\mathbf{G}}$ and its Nyström approximation ($\widehat{\mathbf{G}}$ instead of $\mathbf{G}$ in (3)). The error bound in Lemma II.1 indicates that the approximation error decreases when the singular values of $\widetilde{\mathbf{G}}$ decay rapidly. In particular, if the number of landmark points $m$ approaches the rank of $\widetilde{\mathbf{G}}$, then $\sigma_{m+1}(\widetilde{\mathbf{G}})\to 0$, leading to tighter approximations. The following result formalizes this by guaranteeing perfect reconstruction when the number of landmark points equals the rank of the extended kernel matrix.
Corollary II.2 (Exact recovery)
We next empirically validate the low-rank property of the kernel matrix in pre-trained vision models, which is indeed the assumption underlying Lemma (II.1) and Corollary (II.2). Fig. 1 illustrates the singular value spectrum of the attention matrices from two distinct pre-trained models. In both cases, the rapid singular value decay confirms the low-rank structure of the attention matrices.
(a) SwinIR [2]
(b) Uformer-B [1]
Figure 1: Variation of the top $50$ singular values of attention maps ($N=32^{2}$) from SwinIR and Uformer, averaged over all heads and layers. The steep decay within the first $20$ singular values highlights the low-rank structure of the attention matrices.
III Experiments
@@
We evaluate drop-in replacement capabilities of PnP-Nystra in existing restoration models, followed by an ablation study of its algorithmic parameters. PSNR and SSIM metrics for the experiments are reported, alongside runtime and speed-up on both GPU and CPU platforms. All timings are averaged over multiple inference runs on a single input sample per test dataset. Experiments are conducted using PyTorch $2.7$ with CUDA $12.2$, with GPU and CPU benchmarks performed on NVIDIA RTX $4090$ and Intel Xeon Gold processors, respectively. The implementation is available at: https://github.com/Srinivas-512/PnP_Nystra.
III-A PnP-Nystra for restoration
In this section, we evaluate PnP-Nystra as a drop-in replacement for self-attention layers in three state-of-the-art restoration models: SwinIR [2], Uformer-B [1], and RVRT [4]. All projection weights ($\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$) are retained (no fine-tuning performed) in the pre-trained models. Landmark count $m\in\{16,32\}$ and pseudoinverse iterations ($2$-$6$) are selected to optimize performance across all experiments. We emphasize that our evaluations are restricted to the original pre-trained models, which already deliver state-of-the-art restoration performance. Our goal is not exhaustive benchmarking for restoration but to assess drop-in approximation within high-performing baselines.
III-A1 Image denoising
Figure 2: Comparison of three image patches from Uformer-B and its PnP-Nystra variant for image denoising.
Figure 3: 2×2$\times$2 × super-resolution with SwinIR: PnP-Nystra closely matches the original pre-trained model in PSNR (in dB), SSIM, and visual quality. Error maps (scaled to [0-255]) show pixel-level differences from ground-truth, confirming visual fidelity.
@@ III-A1 Image denoising
TABLE II: Comparison of Uformer-B and Its PnP-Nystra Variant for Image Denoising
Timing Analysis (GPU / CPU, in ms)
Resolution Original PnP-Nystra Speedup
$512$$\times$$512$ $219.33$ / $5760$ $81.96$ / $1580$ 2.67×2.67$\times$2.67 × / 3.65×3.65$\times$3.65 ×
$256$$\times$$256$ $88.77$ / $2310$ $42.41$ / $440$ 2.09×2.09$\times$2.09 × / 5.25×5.25$\times$5.25 ×
Accuracy Analysis (PSNR in dB/ SSIM)
Dataset Original PnP-Nystra Drop
SIDD $38.89$ / $0.8950$ $37.73$ / $0.8836$ $1.16$ / $0.0114$
BSDS200 $28.05$ / $0.8067$ $26.81$ / $0.7748$ $1.24$ / $0.0319$
TABLE III: Comparison of SwinIR and Its PnP-Nystra Variant for Image Super-resolution for different scale factors
Timing Analysis (GPU / CPU in ms)
Resolution Scale Original PnP-Nystra Speedup
$288$$\times$$288$ $2$ $557.32$ / $15330$ $240.80$ / $6902$ 2.31×2.31$\times$2.31 × / $2\times2$
$160$$\times$$160$ $4$ $165.33$ / $4380$ $62.64$ / $1460$ 2.64×2.64$\times$2.64 × / 3.00×3.$0$
$96$$\times$$96$ $8$ $60.02$ / $1810$ $28.28$ / $570$ 2.12×2.12$\times$2.12 × / 3.17×3.17$\times$3.17 ×
Accuracy Analysis (PSNR in dB / SSIM)
Dataset Scale Original PnP-Nystra Drop
Set5 $2$ $36.05$ / $0.9450$ $35.30$ / $0.9412$ $0.75$ / $0.0038$
$4$ $30.71$ / $0.8701$ $29.13$ / $0.8449$ $1.58$ / $0.0252$
$8$ $25.60$ / $0.7403$ $24.40$ / $0.6868$ $1.20$ / $0.0535$
BSDS100 $2$ $30.98$ / $0.8922$ $30.27$ / $0.8830$ $0.71$ / $0.0092$
$4$ $26.43$ / $0.7183$ $25.67$ / $0.6946$ $0.76$ / $0.0237$
$8$ $23.61$ / $0.5680$ $23.04$ / $0.5441$ $0.57$ / $0.0239$
III-A2 Image super-resolution

### v2

# Plug-and-Play Linear Attention with Provable Guarantees for Training-Free Image Restoration

@@ Abstract
Index Terms—  Attention, Image restoration, Training-free, Plug-and-Play, Nyström approximation.
1 Introduction
Transformers have recently demonstrated strong performance across a range of low-level vision tasks, including image denoising, deblurring [21], super-resolution [8], and dehazing [15]. These models rely on a self-attention mechanism that captures long-range dependencies by computing pairwise similarity between features corresponding to image patches. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{i}}$ be the input feature map corresponding to $N$ tokens (e.g., pixels or patches), where each token is represented by a $d_{i}$-dimensional feature vector. To compute self-attention [18], the input $\mathbf{X}$ is first projected into query, key, and value representations:
@@ 1 Introduction
However, the attention operation in (1) incurs a quadratic complexity in the number of input tokens $N$, which corresponds to the number of pixels or patches within a subregion of a frame [18]. In particular, the time and storage complexity for computing (1) is $\mathcal{O}(N^{2}d)$ and $\mathcal{O}(N^{2}+Nd)$, respectively, dominated by storing and computing the large attention matrices $\mathbf{A}$ and $\mathbf{S}$. Currently, most vision models employ Multi-Head Self-Attention (MHSA) [18], an extension of self-attention. In MHSA, the embedding dimension $d$ is divided into $h$ attention heads, each of dimension $d_{h}$, such that $d=hd_{h}$. For each head, self-attention outputs are independently computed, concatenated, and linearly projected to form the MHSA output. The time and storage complexity of MHSA is quadratic in $N$, similar to self-attention. This computational bottleneck makes transformers less suitable for real-time or resource-constrained settings, where convolutional networks are still often preferred due to their computational efficiency [10, 14].
Several approaches approximate self-attention via low-rank [20, 3, 23] or sparse [6] representations to reduce quadratic complexity, but these methods often require retraining. Linformer [20] projects keys/values into lower-dimensional subspaces via learned linear maps, achieving linear complexity in $N$. Performer [3] replaces the softmax kernel with random feature approximation, resulting in linear complexity and unbiased attention estimates. Nyströmformer [23] heuristically reconstructs attention using landmark-based sampling. Axial Attention [7] factorizes 2D attention into 1D along spatial axes, reducing cost from $H^{2}W^{2}$ to $HW(H+W)$, where the number of tokens is $N=HW$. FlashAttention [4] and Triton [17] preserve exact attention while optimizing execution through custom fused GPU kernels, but are hardware-specific and difficult to integrate into pre-trained models.
@@
1) Linear attention: Unlike MHSA, which incurs quadratic time and memory complexity in the number of tokens $N$, PnP-Nystra achieves linear scaling in both time and memory.
2) Impressive acceleration: PnP-Nystra serves as a drop-in replacement for standard self-attention in existing window-attention transformer backbones for image restoration, enabling substantial inference-time acceleration without modifying the pretrained weights. In our experiments, the GPU inference speedup ranges from $1.8\times$ to $3.6\times$ across restoration tasks, while the gains are even larger on CPU inference.
3) Robust performance without retraining: Despite significant acceleration, the reduction in restoration quality is minimal. For instance, in image denoising, PnP-Nystra achieves more than $3.5\times$ speedup on CPU with <1<1 dB drop in PSNR and <0.03<0.03 drop in SSIM.
To the best of our knowledge, we are the first to show that a purely drop-in linear-attention replacement can remain close to the original pretrained self-attention outputs without any finetuning, while retaining provable approximation guarantees.
2 Proposed Method
@@
Note that, while the Nyströmformer [23] also employs the Nyström method, it operates on the softmax-transformed attention matrix $\mathbf{S}$. However, the softmax operation prevents $\mathbf{S}$ from being naturally partitioned into submatrices - a core requirement for Nyström-based approximation. As a result, the authors resort to a heuristic strategy, which makes their approximation devoid of guarantees such as those in Lemma 2.1.
2.1 Approximation error
Lemma 2.1 (Spectral norm error bound).
Assume $\mathbf{G}_{A}$ is nonsingular. Then,
[…]
The proof follows from Lemma 5.10 in [1]. The error bound in Lemma 2.1 indicates that the approximation error decreases when the singular values of $\widetilde{\mathbf{G}}$ decay rapidly. In particular, if the number of landmark points $m$ approaches the rank of $\widetilde{\mathbf{G}}$, then $\sigma_{m+1}(\widetilde{\mathbf{G}})\to 0$, leading to tighter approximations. The following result formalizes this by guaranteeing perfect reconstruction when the number of landmark points equals the rank of the extended kernel matrix.
Corollary 2.2 (Exact recovery).
We next empirically validate the low-rank property of the kernel matrix in pre-trained vision models, which is indeed the assumption underlying Lemma 2.1 and Corollary 2.2. Fig. 1 illustrates the singular value spectrum of the attention matrices from two distinct pre-trained models. In both cases, the rapid singular value decay confirms the low-rank structure of the attention matrices.
(a) SwinIR [8]
(b) Uformer-B [21]
Fig. 1: Variation of the top $50$ singular values of attention maps ($N=32^{2}$) from SwinIR and Uformer, averaged over all heads and layers. The steep decay within the first 2020 singular values highlights the low-rank structure of the attention matrices.
2.2 Computational Details
Standard self-attention has time complexity $\mathcal{O}(N^{2}d)$ and memory $\mathcal{O}(N^{2}+Nd)$. By avoiding explicit $N\times N$ attention, PnP-Nystra reduces the overall complexity to $\mathcal{O}(Nm(d+d_{v}))$ time and $\mathcal{O}(m^{2}+N(m+d_{v}))$ memory for $m\ll N$ (Table 1).
Table 1: Per-step time/space complexity of PnP-Nystra.
Step 2 3 4 5 6 7
@@ 2.2 Computational Details
We form landmarks by average pooling over non-overlapping windows as in [9]. For numerical stability, we replace PyTorch’s pseudoinverse with the iterative method of [13]. Steps 2–3 evaluate exponential kernels, which may overflow and destabilize the Nyström computation. We adopt a softmax-style row-max stabilization by introducing diagonal matrices $\mathbf{M}_{L}\in\mathbb{R}^{N\times N}$ and $\mathbf{M}_{U}\in\mathbb{R}^{m\times m}$ as
$(\mathbf{M}_{L})_{ii}$ $=\exp(m_{L}(i)),$ $\quad m_{L}(i)$ $=\max_{j}\Big(\frac{\mathbf{Q}\bar{\mathbf{K}}^{\top}}{\sqrt{d}}\Big)_{ij}.$
$(\mathbf{M}_{U})_{ii}$ $=\exp(m_{U}(i)),$ $\quad m_{U}(i)$ $=\max_{j}\Big(\frac{\bar{\mathbf{Q}}\mathbf{K}^{\top}}{\sqrt{d}}\Big)_{ij}.$
This rescales each row of $\mathbf{G}_{L}$ and $\mathbf{G}_{U}$ by its maximum log-score. Using $\mathbf{M}_{L}\mathbf{M}_{L}^{-1}=\mathbf{I}$ and $\mathbf{M}_{U}\mathbf{M}_{U}^{-1}=\mathbf{I}$, we rewrite
$\widehat{\mathbf{G}}=\mathbf{M}_{L}\mathbf{M}_{L}^{-1}\mathbf{G}_{L}\,\mathbf{G}_{A}^{\dagger}\,\mathbf{M}_{U}\mathbf{M}_{U}^{-1}\mathbf{G}_{U}=\mathbf{M}_{L}\,\mathbf{G}_{L}^{n}\,(\mathbf{G}_{A}^{n})^{\dagger}\,\mathbf{G}_{U}^{n},$
with normalized blocks
$\mathbf{G}_{L}^{n}:=\mathbf{M}_{L}^{-1}\mathbf{G}_{L},\quad\mathbf{G}_{A}^{n}:=\mathbf{M}_{U}^{-1}\mathbf{G}_{A},\quad\mathbf{G}_{U}^{n}:=\mathbf{M}_{U}^{-1}\mathbf{G}_{U}.$
Equivalently, $\widehat{\mathbf{G}}=\mathbf{M}_{L}\widehat{\mathbf{G}}^{n}$, where $\widehat{\mathbf{G}}^{n}:=\mathbf{G}_{L}^{n}(\mathbf{G}_{A}^{n})^{\dagger}\mathbf{G}_{U}^{n}$ , and the stabilized output in (7) is given as
$\widehat{\mathbf{O}}=(\mathbf{M}_{L}\widehat{\mathbf{G}}^{n}\mathbf{V})\oslash(\mathbf{M}_{L}\widehat{\mathbf{G}}^{n}\mathbf{1}_{N})=(\widehat{\mathbf{G}}^{n}\mathbf{V})\oslash(\widehat{\mathbf{G}}^{n}\mathbf{1}_{N}).$ (8)
In practice, Algorithm 1 computes $\widehat{\mathbf{O}}$ using the stabilized form in (8) instead of (7).
Fig. 2: Single-image dehazing. Top: hazy input and restored outputs. Bottom: ground truth and pixel-wise error maps (scaled to [0–255]). PnP-Nystra matches the pretrained model closely while producing cleaner dehazing than other training-free linear attention baselines, as evidenced by metrics (PSNR(dB), SSIM) and reconstruction of trees and houses in the zoomed area.
3 Experiments
@@
We evaluate drop-in replacement capabilities of PnP-Nystra in existing restoration models, followed by an ablation study of its algorithmic parameters. PSNR and SSIM metrics for the experiments are reported, alongside runtime and speedup on both GPU and CPU platforms. All timings are averaged over multiple inference runs per test dataset. Experiments are conducted using PyTorch $2.7$ with CUDA $12.2$, with GPU and CPU benchmarks performed on NVIDIA RTX $4090$ and Intel Xeon Gold processors, respectively. The implementation is available at: https://anonymous.4open.science/r/PnP_Nystra-0F31.
3.1 PnP-Nystra for restoration
In this section, we evaluate PnP-Nystra as a drop-in replacement for MHSA in three state-of-the-art restoration Transformers: SwinIR [8], Uformer-B [21], and Dehazeformer [15]. We keep all pretrained projection weights $(\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V})$ fixed and perform no finetuning. Unless stated otherwise, we use $m=16$ landmarks and 4466 pseudoinverse iterations, chosen to provide strong performance across tasks. We compare against other training-free linear attention substitutes, such as Performer [3] and Nyströmformer [23]. Our intent is not exhaustive restoration benchmarking, but to quantify the inference-time trade-off of attention approximation within these high-performing pretrained baselines. Speedups over the original attention are summarized in Table 2.
Table 2: Timing comparison per attention block for original model and PnP-Nystra. GPU/CPU time (in ms) is reported.
Resolution Scale Original PnP-Nystra Speedup
SwinIR for Super-Resolution
$88\times 288 22$ $524.61$ / $17916$ $235.76$ / $8139$ $23\times / 2.2$$20\times$
$160\times 160$} 4{4} $159.96$ / $5298$ $65.46$ / $1688$ $2.44\times$ / $3.13\times$
$96\times 96$ 88 $59.18$ / $2274$ $29.11$ / $652$ $2.03\times$ / $3.49\times$
Uformer-B for Denoising
$512\times 512$ $218.01$ / $6958$ $67.26$ / $1853$ $3.24\times$ / $3.75\times$
$256\times 256$ $81.81$ / $2610$ $22.64$ / $372$ $3.61\times$ / $7.05\times$
Uformer-B for Deblurring
$768\times 768$} $285.46$ / $9098$ $155.19$ / $4942$ $\times / 1.84×1.84\times$
Dehazeformer for Dehazing
$416\times 576$ $143.92$ / $5106$ $51.71$ / $1689$ $2.78\times$ / $3.02\times$
$576\times 576$ $183.53$ / $6938$ $70.05$ / $2558$ $2.62\times$ / $2.71\times$
$736\times 576$ $261.80$ / $10192$ $105.02$ / $4157$ $2.49\times$ / $2.45\times$
3.1.1 Dehazing
We evaluate Dehazeformer [15] with PnP-Nystra as a drop-in replacement for MHSA on single-image dehazing. We use a window size of $32$ (i.e., $N=32^{2}$ tokens) and run $5$ pseudoinverse iterations for all dehazing experiments. Table 3 reports PSNR/SSIM on different datasets, where PnP-Nystra consistently outperforms other training-free linear attention substitutes, producing accuracy close to (and in some cases exceeding) the original pretrained Dehazeformer. This is achieved at a $2.5\times$ to $2.8\times$ speedup as shown in the Table. 2. Qualitative results in Fig. 2 show that PnP-Nystra preserves fine details and contrast on par with the original model, while producing cleaner haze removal than other training-free approximations.
Table 3: Accuracy Comparison for Image Dehazing (PSNR / SSIM) with other linear attention methods
Dataset Original Performer Nystromformer PnP-Nystra
Reside-Out $26.51$ / $0.9512$ $25.26$ / $0.9328$ $23.57$ / $0.9266$ 26.34 / 0.9510
RSHaze $32.80$ / $0.9261$ $29.00$ / $0.8957$ $33.18$ / $0.9363$ 33.58 / 0.9405
RESIDE-6K $24.79$ / $0.9352$ $21.00$ / $0.8834$ $21.37$ / $0.8962$ 23.08 / 0.9163
3.1.2 Denoising and Deblurring
@@
Table 4: Accuracy comparison (PSNR / SSIM) for Uformer-B with training-free linear attention substitutes.
Dataset Original Performer Nyströmformer PnP-Nystra
Denoising
SIDD $38.89$ / $0.8950$ $34.80$ / $0.8713$ $37.46$ / $0.8838$ $\mathbf{37.88}$ / $\mathbf{0.8842}$
BSDS200 $28.05$ / $0.8067$ $24.80$ / $0.7469$ $26.77$ / $0.7829$ $\mathbf{27.17}$ / $\mathbf{0.7838}$
Deblurring
RealBlur-R $33.98$ / $0.9404$ $33.97$ / $0.9384$ $34.01$ / $0.9407$ $\mathbf{34.02}$ / $\mathbf{0.9407}$
3.1.3 Super-resolution